100 cm of brass increased by 0.18 cm when its temperature rises by 95°C. What is the linear expansivity of brass?


1
Expert's answer
2022-02-18T15:53:42-0500

Given:

"l_0=100\:\rm cm"

"\Delta l=0.18\:\rm cm"

"\Delta t=95\:\rm K"


The thermal expansion is described by an equation


"\Delta l=l_0\alpha\Delta t"

Hence, the linear expansivity


"\alpha=\Delta l/(l_0\Delta t)""\alpha=\rm 0.18\: cm/(100\: cm*95 \: K)=1.9*10^{-5}\: K^{-1}"